\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Kashaev:1996rz}
R.~M. Kashaev and S.~M. Sergeev, ``{On pentagon, ten term, and tetrahedron
  relations},'' \href{http://arxiv.org/abs/q-alg/9607032}{{\ttfamily
  arXiv:q-alg/9607032}}.

\bibitem{Kashaev:2014rea}
R.~M. Kashaev, ``{Beta pentagon relations},''
  \href{http://dx.doi.org/10.1007/s11232-014-0208-4}{{\em Theor. Math. Phys.}
  {\bfseries 181} no.~1, (2014) 1194--1205},
  \href{http://arxiv.org/abs/1403.1298}{{\ttfamily arXiv:1403.1298 [math-ph]}}.

\bibitem{Bozkurt:2018xno}
D.~N. Bozkurt and I.~Gahramanov, ``{Pentagon identities arising in
  supersymmetric gauge theory computations},''
  \href{http://dx.doi.org/10.1134/S0040577919020028}{{\em Teor. Mat. Fiz.}
  {\bfseries 198} no.~2, (2019) 215--224},
\href{http://arxiv.org/abs/1803.00855}{{\ttfamily arXiv:1803.00855 [math-ph]}}.
%%CITATION = ARXIV:1803.00855;%%.

\bibitem{Bozkurt:2020gyy}
D.~N. Bozkurt, I.~Gahramanov, and M.~Mullahasanoglu, ``{Lens partition
  function, pentagon identity, and star-triangle relation},''
  \href{http://dx.doi.org/10.1103/PhysRevD.103.126013}{{\em Phys. Rev. D}
  {\bfseries 103} no.~12, (2021) 126013},
  \href{http://arxiv.org/abs/2009.14198}{{\ttfamily arXiv:2009.14198
  [hep-th]}}.

\bibitem{Allman:2017xma}
J.~Allman and R.~Rim\'anyi, ``{Quantum dilogarithm identities for the square
  product of A-type Dynkin quivers},''
  \href{http://dx.doi.org/10.4310/MRL.2018.v25.n4.a1}{{\em Math. Res. Lett.}
  {\bfseries 25} (2018) 1037--1087},
  \href{http://arxiv.org/abs/1702.04766}{{\ttfamily arXiv:1702.04766
  [math.RT]}}.

\bibitem{Gahramanov:2014ona}
I.~Gahramanov and H.~Rosengren, ``{Integral pentagon relations for 3d
  superconformal indices},'' \href{http://arxiv.org/abs/1412.2926}{{\ttfamily
  arXiv:1412.2926 [hep-th]}}.
[Proc. Symp. Pure Math.93,165(2016)].
%%CITATION = ARXIV:1412.2926;%%.

\bibitem{Gahramanov:2013rda}
I.~Gahramanov and H.~Rosengren, ``{A new pentagon identity for the tetrahedron
  index},'' \href{http://dx.doi.org/10.1007/JHEP11(2013)128}{{\em JHEP}
  {\bfseries 11} (2013) 128},
\href{http://arxiv.org/abs/1309.2195}{{\ttfamily arXiv:1309.2195 [hep-th]}}.
%%CITATION = ARXIV:1309.2195;%%.

\bibitem{Gahramanov:2016wxi}
I.~Gahramanov and H.~Rosengren, ``{Basic hypergeometry of supersymmetric
  dualities},'' \href{http://dx.doi.org/10.1016/j.nuclphysb.2016.10.004}{{\em
  Nucl. Phys.} {\bfseries B913} (2016) 747--768},
\href{http://arxiv.org/abs/1606.08185}{{\ttfamily arXiv:1606.08185 [hep-th]}}.
%%CITATION = ARXIV:1606.08185;%%.

\bibitem{Gahramanov:2022jxz}
I.~Gahramanov, B.~Keskin, D.~Kosva, and M.~Mullahasanoglu, ``{On Bailey pairs
  for $\mathcal N=2$ supersymmetric gauge theories on $S_b^3/\mathbb{Z}_r$},''
  \href{http://arxiv.org/abs/2210.11455}{{\ttfamily arXiv:2210.11455
  [hep-th]}}.

\bibitem{Gahramanov:2021pgu}
I.~Gahramanov and O.~E. Kaluc, ``{Bailey pairs for the q-hypergeometric
  integral pentagon identity},''
  \href{http://arxiv.org/abs/2111.14793}{{\ttfamily arXiv:2111.14793
  [math-ph]}}.

\bibitem{Dede:2022ofo}
M.~Dede, ``{A comment on the solutions of the generalized Faddeev-Volkov
  model},'' \href{http://arxiv.org/abs/2206.14271}{{\ttfamily arXiv:2206.14271
  [math-ph]}}.

\bibitem{Kashaev:2012cz}
R.~Kashaev, F.~Luo, and G.~Vartanov, ``{A TQFT of Turaev\textendash{}Viro Type
  on Shaped Triangulations},''
  \href{http://dx.doi.org/10.1007/s00023-015-0427-8}{{\em Annales Henri
  Poincare} {\bfseries 17} no.~5, (2016) 1109--1143},
  \href{http://arxiv.org/abs/1210.8393}{{\ttfamily arXiv:1210.8393 [math.QA]}}.

\bibitem{colazzo2020set}
I.~Colazzo, E.~Jespers, and {\L}.~Kubat, ``Set-theoretic solutions of the
  pentagon equation,'' {\em Communications in Mathematical Physics} {\bfseries
  380} no.~2, (2020) 1003--1024.

\bibitem{kashaev1997heisenberg}
R.~Kashaev, ``Heisenberg double and the pentagon relation,'' in {\em St.
  Petersburg Math. J}, Citeseer.
\newblock 1997.

\bibitem{militaru2004heisenberg}
G.~Militaru, ``Heisenberg double, pentagon equation, structure and
  classification of finite-dimensional hopf algebras,'' {\em Journal of the
  London Mathematical Society} {\bfseries 69} no.~1, (2004) 44--64.

\bibitem{aghaei2019heisenberg}
N.~Aghaei and M.~Pawelkiewicz, ``Heisenberg double and drinfeld double of the
  quantum superplane,'' {\em arXiv preprint arXiv:1909.04565} (2019) .

\bibitem{Imamura:2012rq}
Y.~Imamura and D.~Yokoyama, ``{$S^3/Z_n$ partition function and dualities},''
  \href{http://dx.doi.org/10.1007/JHEP11(2012)122}{{\em JHEP} {\bfseries 11}
  (2012) 122},
\href{http://arxiv.org/abs/1208.1404}{{\ttfamily arXiv:1208.1404 [hep-th]}}.
%%CITATION = ARXIV:1208.1404;%%.

\bibitem{Imamura:2013qxa}
Y.~Imamura, H.~Matsuno, and D.~Yokoyama, ``{Factorization of the
  $S^3/\mathbb{Z}_n$ partition function},''
  \href{http://dx.doi.org/10.1103/PhysRevD.89.085003}{{\em Phys. Rev.}
  {\bfseries D89} no.~8, (2014) 085003},
\href{http://arxiv.org/abs/1311.2371}{{\ttfamily arXiv:1311.2371 [hep-th]}}.
%%CITATION = ARXIV:1311.2371;%%.

\bibitem{Benini:2011nc}
F.~Benini, T.~Nishioka, and M.~Yamazaki, ``{4d Index to 3d Index and 2d
  TQFT},'' \href{http://dx.doi.org/10.1103/PhysRevD.86.065015}{{\em Phys. Rev.}
  {\bfseries D86} (2012) 065015},
\href{http://arxiv.org/abs/1109.0283}{{\ttfamily arXiv:1109.0283 [hep-th]}}.
%%CITATION = ARXIV:1109.0283;%%.

\bibitem{Yamazaki:2013fva}
M.~Yamazaki, ``{Four-dimensional superconformal index reloaded},''
  \href{http://dx.doi.org/10.1007/s11232-013-0012-6}{{\em Theor. Math. Phys.}
  {\bfseries 174} (2013) 154--166}.
[Teor. Mat. Fiz.174,177(2013)].
%%CITATION = TMPHA,174,154;%%.

\bibitem{Eren:2019ibl}
E.~Eren, I.~Gahramanov, S.~Jafarzade, and G.~Mogol, ``{Gamma function solutions
  to the star-triangle equation},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2020.115283}{{\em Nucl. Phys. B}
  {\bfseries 963} (2021) 115283},
  \href{http://arxiv.org/abs/1912.12271}{{\ttfamily arXiv:1912.12271
  [math-ph]}}.

\bibitem{Gahramanov:2016ilb}
I.~Gahramanov and A.~P. Kels, ``{The star-triangle relation, lens partition
  function, and hypergeometric sum/integrals},''
  \href{http://dx.doi.org/10.1007/JHEP02(2017)040}{{\em JHEP} {\bfseries 02}
  (2017) 040},
\href{http://arxiv.org/abs/1610.09229}{{\ttfamily arXiv:1610.09229 [math-ph]}}.
%%CITATION = ARXIV:1610.09229;%%.

\bibitem{Catak:2021coz}
E.~Catak, I.~Gahramanov, and M.~Mullahasanoglu, ``{Hyperbolic and trigonometric
  hypergeometric solutions to the star-star equation},''
  \href{http://dx.doi.org/10.1140/epjc/s10052-022-10661-x}{{\em Eur. Phys. J.
  C} {\bfseries 82} no.~9, (2022) 789},
  \href{http://arxiv.org/abs/2107.06880}{{\ttfamily arXiv:2107.06880
  [hep-th]}}.

\bibitem{Nieri:2015yia}
F.~Nieri and S.~Pasquetti, ``{Factorisation and holomorphic blocks in 4d},''
  \href{http://dx.doi.org/10.1007/JHEP11(2015)155}{{\em JHEP} {\bfseries 11}
  (2015) 155},
\href{http://arxiv.org/abs/1507.00261}{{\ttfamily arXiv:1507.00261 [hep-th]}}.
%%CITATION = ARXIV:1507.00261;%%.

\bibitem{Sarkissian:2018ppc}
G.~Sarkissian and V.~P. Spiridonov, ``{From rarefied elliptic beta integral to
  parafermionic star-triangle relation},''
  \href{http://dx.doi.org/10.1007/JHEP10(2018)097}{{\em JHEP} {\bfseries 10}
  (2018) 097},
\href{http://arxiv.org/abs/1809.00493}{{\ttfamily arXiv:1809.00493 [hep-th]}}.
%%CITATION = ARXIV:1809.00493;%%.

\bibitem{Mullahasanoglu:2021xyf}
M.~Mullahasanoglu and N.~Tas, ``{Lens Partition Functions and Integrability
  Properties},'' \href{http://arxiv.org/abs/2112.15161}{{\ttfamily
  arXiv:2112.15161 [hep-th]}}.

\bibitem{Gahramanov:2015cva}
I.~Gahramanov and V.~P. Spiridonov, ``{The star-triangle relation and 3d
  superconformal indices},''
  \href{http://dx.doi.org/10.1007/JHEP08(2015)040}{{\em JHEP} {\bfseries 08}
  (2015) 040},
\href{http://arxiv.org/abs/1505.00765}{{\ttfamily arXiv:1505.00765 [hep-th]}}.
%%CITATION = ARXIV:1505.00765;%%.

\bibitem{Spiridonov:2010em}
V.~P. Spiridonov, ``{Elliptic beta integrals and solvable models of statistical
  mechanics},'' {\em Contemp. Math.} {\bfseries 563} (2012) 181--211,
\href{http://arxiv.org/abs/1011.3798}{{\ttfamily arXiv:1011.3798 [hep-th]}}.
%%CITATION = ARXIV:1011.3798;%%.

\bibitem{Kels:2013ola}
A.~P. Kels, ``{A new solution of the star-triangle relation},''
  \href{http://dx.doi.org/10.1088/1751-8113/47/5/055203}{{\em J. Phys.}
  {\bfseries A47} (2014) 055203},
\href{http://arxiv.org/abs/1302.3025}{{\ttfamily arXiv:1302.3025 [math-ph]}}.
%%CITATION = ARXIV:1302.3025;%%.

\bibitem{Intriligator:1996ex}
K.~A. Intriligator and N.~Seiberg, ``{Mirror symmetry in three-dimensional
  gauge theories},'' \href{http://dx.doi.org/10.1016/0370-2693(96)01088-X}{{\em
  Phys. Lett.} {\bfseries B387} (1996) 513--519},
\href{http://arxiv.org/abs/hep-th/9607207}{{\ttfamily arXiv:hep-th/9607207
  [hep-th]}}.
%%CITATION = HEP-TH/9607207;%%.

\bibitem{Aharony:1997bx}
O.~Aharony, A.~Hanany, K.~A. Intriligator, N.~Seiberg, and M.~Strassler,
  ``{Aspects of N=2 supersymmetric gauge theories in three-dimensions},''
  \href{http://dx.doi.org/10.1016/S0550-3213(97)00323-4}{{\em Nucl.Phys.}
  {\bfseries B499} (1997) 67--99},
\href{http://arxiv.org/abs/hep-th/9703110}{{\ttfamily arXiv:hep-th/9703110
  [hep-th]}}.
%%CITATION = HEP-TH/9703110;%%.

\bibitem{Sarkissian:2022mlb}
G.~Sarkissian and V.~Spiridonov, ``{Elliptic, hyperbolic, complex gamma
  functions and QFT in various dimensions},''
  \href{http://dx.doi.org/10.22323/1.412.0037}{{\em PoS} {\bfseries Regio2021}
  (2022) 037}.

\bibitem{Faddeev:2000if}
L.~D. Faddeev, R.~M. Kashaev, and A.~Y. Volkov, ``{Strongly coupled quantum
  discrete Liouville theory. 1. Algebraic approach and duality},''
  \href{http://dx.doi.org/10.1007/s002200100412}{{\em Commun. Math. Phys.}
  {\bfseries 219} (2001) 199--219},
  \href{http://arxiv.org/abs/hep-th/0006156}{{\ttfamily arXiv:hep-th/0006156}}.

\bibitem{Dimofte:2011ju}
T.~Dimofte, D.~Gaiotto, and S.~Gukov, ``{Gauge Theories Labelled by
  Three-Manifolds},'' \href{http://dx.doi.org/10.1007/s00220-013-1863-2}{{\em
  Commun. Math. Phys.} {\bfseries 325} (2014) 367--419},
  \href{http://arxiv.org/abs/1108.4389}{{\ttfamily arXiv:1108.4389 [hep-th]}}.

\bibitem{Dimofte:2012pd}
T.~Dimofte and D.~Gaiotto, ``{An E7 Surprise},''
  \href{http://dx.doi.org/10.1007/JHEP10(2012)129}{{\em JHEP} {\bfseries 10}
  (2012) 129}, \href{http://arxiv.org/abs/1209.1404}{{\ttfamily arXiv:1209.1404
  [hep-th]}}.

\bibitem{suzuki2002turaev}
K.~Suzuki and M.~Wakui, ``On the turaev-viro-ocneanu invariant of 3-manifolds
  derived from the $e_6$-subfactor,'' {\em Kyushu Journal of Mathematics}
  {\bfseries 56} no.~1, (2002) 59--81.

\bibitem{Gahramanov:2017ysd}
I.~Gahramanov and S.~Jafarzade, ``{Integrable lattice spin models from
  supersymmetric dualities},''
  \href{http://dx.doi.org/10.1134/S1547477118060079}{{\em Phys. Part. Nucl.
  Lett.} {\bfseries 15} no.~6, (2018) 650--667},
\href{http://arxiv.org/abs/1712.09651}{{\ttfamily arXiv:1712.09651 [math-ph]}}.
%%CITATION = ARXIV:1712.09651;%%.

\bibitem{Yamazaki:2018xbx}
M.~Yamazaki, ``{Integrability As Duality: The Gauge/YBE Correspondence},''
  \href{http://dx.doi.org/10.1016/j.physrep.2020.01.006}{{\em Phys. Rept.}
  {\bfseries 859} (2020) 1--20},
  \href{http://arxiv.org/abs/1808.04374}{{\ttfamily arXiv:1808.04374
  [hep-th]}}.

\bibitem{Alexandrov:2015xir}
S.~Alexandrov and B.~Pioline, ``{Theta series, wall-crossing and quantum
  dilogarithm identities},''
  \href{http://dx.doi.org/10.1007/s11005-016-0857-3}{{\em Lett. Math. Phys.}
  {\bfseries 106} no.~8, (2016) 1037--1066},
  \href{http://arxiv.org/abs/1511.02892}{{\ttfamily arXiv:1511.02892
  [hep-th]}}.

\bibitem{Faddeev:2011tah}
L.~D. Faddeev, ``{Volkov's Pentagon for the Modular Quantum Dilogarithm},''
  \href{http://dx.doi.org/10.1007/s10688-011-0031-8}{{\em Funct. Anal. Appl.}
  {\bfseries 45} (2011) 291}, \href{http://arxiv.org/abs/1201.6464}{{\ttfamily
  arXiv:1201.6464 [math.QA]}}.

\end{thebibliography}\endgroup

\documentclass[11pt,a4paper]{article}
\usepackage{jheppub}
\usepackage{amsmath}
\usepackage[most]{tcolorbox}
\usepackage{dsfont}
\usepackage{ulem}
\usepackage{natbib}
\usepackage{xcolor}
\usepackage[hang,flushmargin]{footmisc}
\usepackage{tikz-cd}
\usepackage{enumitem}
\usepackage{amsfonts,amssymb, amscd,amsmath,latexsym,amsbsy,bm}
\usepackage{stmaryrd}
\usepackage{todonotes}
\usepackage{stackrel}
\usepackage{amsthm}

\usepackage{graphicx}

\title{Notes on the lens integral pentagon identity}

\author{H. Kubra Bag$^{a}$, Osman Ergec$^{a}$, and Ilmar Gahramanov$^{a,b,c}$}
\affiliation{
	$^a$ {Department of Physics, Bogazici University, 34342 Bebek, Istanbul, Turkey}\\[-0.5cm]
	
	$^{b}$ Steklov Mathematical Institute of Russ. Acad. Sci., Gubkina str. 8, 119991 Moscow, Russia\\[-0.5cm]
	
	$^{c}$ Department of Mathematics, Khazar University,  Mehseti St. 41, AZ1096, Baku, Azerbaijan \\
}
\vspace{0.2cm}

\emailAdd{hatice.bag@boun.edu.tr}
\emailAdd{osman.ergec@boun.edu.tr}
\emailAdd{ilmar.gahramanov@boun.edu.tr}

\abstract{We obtain the lens integral pentagon identity for three-dimensional mirror dual theories in terms of hyperbolic hypergeometric functions via reduction of equality for $S_b^3/\mathbb Z_r$ $\mathcal N=2$ supersymmetric partition functions of a certain supersymmetric IR duality. }

\keywords{hyperbolic hypergeometric function, pentagon
	identity, supersymmetric duality}

\begin{document}
	\maketitle
	\flushbottom

\section{Introduction}

Pentagon relations have plenty of important applications in the theory of integrable lattice models of statistical mechanics, topological field theories, knot invariants, etc  \cite{Kashaev:1996rz,Kashaev:2014rea,Bozkurt:2018xno,Bozkurt:2020gyy,Allman:2017xma,Gahramanov:2014ona,Gahramanov:2013rda,Gahramanov:2016wxi,Gahramanov:2022jxz,Gahramanov:2021pgu,Dede:2022ofo,Kashaev:2012cz,colazzo2020set}. From an algebraic perspective, pentagon relations are related to the Heisenberg double \cite{kashaev1997heisenberg,militaru2004heisenberg,aghaei2019heisenberg}. In this paper, we obtain an integral pentagon relation arising from supersymmetric gauge theory computations. In order to obtain new relation, we consider a certain three-dimensional $\mathcal N=2$ supersymmetric duality on lens space $S_b^3/\mathbb Z_r$. By taking several limits of the partition functions for supersymmetric theories we end up with a new pentagon identity for hyperbolic hypergeometric functions.

\section{Notations}

In this paper we use the so-called hyperbolic gamma function which has the following integral representation:
\begin{equation}
    \gamma ^{(2)}(y; \omega_1 , \omega_2) = \exp{\bigg[ - \int_{0}^{\infty} \bigg( \frac{\sinh{(2y - \omega_1 - \omega_2)}x}{2\sinh{(\omega_1 x)}\sinh{(\omega_2 x)}} - \frac{2y-\omega_1 - \omega_2}{2\omega_1\omega_2 x} \bigg) \frac{dx}{x} \bigg]},
\end{equation}
The $\gamma^{(2)}(y;\omega_1, \omega_2)$ function has following asymptotics:
\begin{equation}
    \begin{split}
    \lim_{y\to \infty} e^{\frac{i\pi}{2}B_{2,2}(y;\omega_1,\omega_2)}\gamma^{(2)}(y;\omega_1,\omega_2)=1, \arg{\omega_1} <\arg{y} < \arg{\omega_2} + \pi,
    \\
    \lim_{y\to \infty} e^{-\frac{i\pi}{2}B_{2,2}(y;\omega_1,\omega_2)}\gamma^{(2)}(y;\omega_1,\omega_2)=1, \arg{\omega_1} - \pi <\arg{y} < \arg{\omega_2},
    \end{split}\label{eq:asymptotics}
\end{equation}
where the second order Bernoulli polynomial $B_{2,2}(y;\omega_1,\omega_2)$ is defined as follows
\begin{equation}
    B_{2,2}(y;\omega_1,\omega_2) = \frac{y^2}{\omega_1\omega_2} - \frac{y}{\omega_1} - \frac{y}{\omega_2} + \frac{1}{6}\bigg(\frac{\omega_1}{\omega_2} + \frac{\omega_2}{\omega_1} \bigg) +\frac{1}{2} .
\end{equation}

\section{Lens partition function}

In \cite{Imamura:2012rq,Imamura:2013qxa}, the lens partition function was derived via the supersymmetric localization technique, and in \cite{Benini:2011nc,Yamazaki:2013fva,Eren:2019ibl,Gahramanov:2016ilb}, authors used the dimensional reduction of the four-dimensional lens superconformal index. Here we briefly outline some basic ingredients of $\mathcal N=2$ $S_b^3/\mathbb Z_r$ supersymmetric partition function. We mostly follow the notations of \cite{Catak:2021coz,Bozkurt:2020gyy}. A reader interested in details is encouraged to look at \cite{Imamura:2012rq,Imamura:2013qxa,Nieri:2015yia,Gahramanov:2016ilb}. 

By making the identification $(x,y)\sim(e^{\frac{2\pi i}{r}}x,e^{\frac{2\pi i}{r}}y)$ one can obtain the lens space $S_b^3/\mathbb Z_r$ from the squashed three sphere 
\begin{equation}
    S_b^3= \{(x,y)\in\mathbb{C}^2, b^2|x|^2+b^{-2}|y|^2=1\} \;.
\end{equation}
The supersymmetric partition function on this manifold can be reduced to the following matrix model via supersymmetric localization
\begin{equation}
    Z=\sum_m\int \frac{1}{|W|}\prod_j^{rank G} \frac{dz_j}{2\pi i r}Z_{\text{classical}}[z,m]Z_{\text{one-loop}}[z,m] \;,
\end{equation}
where the sum is over the holonomies $m=\frac{r}{2\pi}\int_C A_\mu dx^\mu$, $C$ is the non-trivial cycle on $S^3_b/\mathbb{Z}_r$ and $A_\mu$ is the gauge field, and the integral is over the Cartan subalgebra of the gauge group and $z_j$ variables stand for the gauge fugacities. The prefactor $|W|$ corresponds to the order of the gauge group, which is broken by the holonomy into a product of $r$ subgroups. 

The classical term $Z_{\text{classical}}$ includes the contributions coming from the classical action of the Chern-Simons term and Fayet-Iliopoulos term.

The one-loop contribution of chiral multiplets is given in terms of hyperbolic hypergeometric function. A chiral multiplet in the fundamental representation of the gauge group contributes as 
\begin{equation}
    \begin{split}
        Z_{\rm chiral} =
\prod_{\rho}\prod_{\phi} & e^{\frac{i\pi}{2r} ((\rho(m) + \phi(n)) (r-\rho(m) - \phi(n))  - (r-1)(\rho(m)+\phi(n))^2 )} \\
& \times\gamma^2(i(\rho(z) + \phi(\Phi)) + \omega_1 (\rho(m) + \phi(n)) + (\omega_1 + \omega_2)/2; \omega_1 r, \omega_1 + \omega_2) \\
& \times \gamma^2(iz + \omega_2(r-(\rho(m) - \phi(n) ) + (\omega_1 + \omega_2)/2 ; \omega_2 r, \omega_1 + \omega_2) .
    \end{split}
\end{equation}
where $\rho$ and $\phi$ weights of the representation of the gauge group, and the flavor group, respectively. Note that the Weyl weight of a chiral multiple is absorbed into the flavor fugacity. The one-loop contribution of the vector multiplet in our case is trivial since we consider only $U(1)$ gauge theories. 

The partition function is completely determined by the group-theoretical data of the theory. Therefore once we know the group-theoretical data of a theory on $S_b^3/\mathbb Z_r$, we can write down the partition function in terms of hyperbolic hypergeometric integral.

\section{Supersymmetric duality}

Our starting point is the following supersymmetric three-dimensional $\mathcal N=2$ IR duality:

\begin{itemize}

\item \textbf{Theory A:} Theory with $U(1)$ gauge symmetry and $SU(3)_L \times SU(3)_R$ flavor group, having three fundamental chiral multiples and three anti-fundamental chiral multiplets and no Chern-Simons term.

\item \textbf{Theory B:} Theory  without gauge degrees of freedom having the same global symmetries. There are  nine ``mesons'', transforming in the fundamental representation of the flavor group $SU(3)_L \times SU(3)_R$.

\end{itemize}

Partition functions of dual theories agree and as a result one obtains the following hyperbolic hypergeometric integral identity \cite{Sarkissian:2018ppc}
\begin{align}
        \sum\limits_{y = 0}^{[r/2]} \epsilon(y)  \int_{-i\infty}^{i\infty} \nonumber
        \prod _{i=1}^{3}\gamma^{(2)}(a_i -z +\omega_1(u_i - y); \omega_1 r, \omega_1 + \omega_2) \gamma^{(2)}(a_i - z +\omega_2(r-(u_i - y)); \omega_2 r, \omega_1 + \omega_2)\\ \nonumber
        \times \gamma^{(2)}( b_i + z + \omega_1(v_i + y); \omega_1 r, \omega_1 + \omega_2) \gamma^{(2)}(b_i + z + \omega_2(r-(v_i + y)); \omega_2 r, \omega_1 + \omega_2) \frac{dz}{ir\sqrt{\omega_1\omega_2}}\\ 
        = \prod\limits_{i,j=1}^{3} \gamma^{(2)}(a_i+ b_j +\omega_1(u_i + v_j); \omega_1 r, \omega_1 + \omega_2) \gamma^{(2)}(a_i + b_j + \omega_2(r-(u_i + v_j)); \omega_2 r, \omega_1 + \omega_2)\label{eq:bailey-limit}
\end{align}
with the balancing conditions $\sum_i a_i+b_i=\omega_1+\omega_2$ and $\sum_i u_i+v_i=r$. Here $a_i$ are $SU(3)_L$ flavor fugacities and $b_i$ are $SU(3)_R$ flavor fugacities. The $\epsilon(y)$ function is defined as $\epsilon(0)=\epsilon(\lfloor\frac{r}{2}\rfloor)=1$ and $\epsilon(y)=2$ otherwise.

Note that this integral identity can be written as the star-triangle relation and pentagon identity \cite{Bozkurt:2020gyy,Mullahasanoglu:2021xyf,Sarkissian:2018ppc}. The Bailey pair for this identity was constructed in \cite{Gahramanov:2022jxz} (see, also \cite{Gahramanov:2015cva}).  A similar identity was discussed for the $S^3$ sphere partition functions (without summation) in \cite{Spiridonov:2010em,Kashaev:2012cz} (see also \cite{Kels:2013ola}) and for the $S^2 \times S^1$ partition function in \cite{Gahramanov:2013rda,Gahramanov:2014ona,Gahramanov:2016wxi}.

Another supersymmetric duality considered in this paper is mirror symmetry. Three-dimensional mirror symmetry was introduced in \cite{Intriligator:1996ex} for $\mathcal N = 4$ supersymmetric gauge theories, and   in \cite{Aharony:1997bx} for $\mathcal N=2$ gauge theories. The simplest example of $\mathcal N = 2$ mirror symmetry is the duality between supersymmetric quantum electrodynamics with one flavor and the free Wess-Zumino theory:

\begin{itemize}

\item \textbf{Theory A:} Theory has one flavor consisting of two chiral fields and one vector multiplet. This theory possesses two U$(1)$ global symmetries: one is the topological U$(1)_{J}$, and the other is the flavor symmetry U$(1)_{A}$. 

\item \textbf{Theory B:} The mirror theory, the free Wess-Zumino model (this theory is often called the XYZ model) is the theory containing three chiral fields interacting through the trilinear superpotential. This theory has also two U$(1)$ global symmetries, named U$(1)_{V}$ and U$(1)_{A}$.

\end{itemize}

In the context of mirror symmetry, one  identifies U$(1)_{J}$ and U$(1)_{A}$ of the SQED theory with U$(1)_{V}$ and U$(1)_{A}$ of the corresponding Wess-Zumino model. The equality of lens partition functions for mirror dual theories can be written as follows
\begin{align} \nonumber
        \sum\limits_{y = 0}^{[r/2]} \epsilon(y)  \int_{-i\infty}^{i\infty} &e^{\frac{2\pi i}{\omega_1 \omega_2 r}(z + \alpha) (\frac{3\alpha + \beta - \omega}{2}) + \frac{i \pi}{r}(y + n_{\alpha}) (2n_{\beta} - r)}
        \gamma^{(2)}(z + \alpha +\omega_1(y+n_{\alpha}); \omega_1 r, \omega)\\ \nonumber
        &\times \gamma^{(2)}(z + \alpha +\omega_2(r-y-n_{\alpha})); \omega_2 r, \omega)\gamma^{(2)}(-z+\alpha +\omega_1(-y + n_{\alpha}); \omega_1 r, \omega)\\ \nonumber
        &\times \gamma^{(2)}(-z+\alpha +\omega_2(r + y -n_{\alpha})); \omega_2 r, \omega)\frac{dz}{ir\sqrt{\omega_1\omega_2}} \\ \nonumber
        &=e^{\frac{2\pi i }{\omega_1 \omega_2 r} (\alpha) (\frac{2\beta - \omega}{2}) + \frac{i \pi}{r} ( -n_{\alpha})(2n_{\beta} - r)} 
        \gamma^{(2)}(\omega - \alpha - \beta + \omega_1(r - n_{\alpha} - n_{\beta}); \omega_1 r, \omega)\\ \nonumber
        &\times \gamma^{(2)}(\omega - \alpha - \beta + \omega_2(n_{\alpha} + n_{\beta});\omega_2 r, \omega)
        \gamma^{(2)}(\beta - \alpha + \omega_1 (n_{\beta} - n_{\alpha}); \omega_1 r, \omega) \\ \nonumber
        & \times \gamma^{(2)}( \beta - \alpha + \omega_2(r-n_{\beta}+n_{\alpha});\omega_2 r, \omega)\\ 
        & \times \gamma^{(2)}(2\alpha + \omega_1 (2n_{\alpha}); \omega_1 r, \omega)\gamma^{(2)}( 2\alpha + \omega_2(r- 2n_{\alpha});\omega_2 r, \omega) \;,  \label{mirror}
\end{align}
where $\alpha$, $\beta$ are fugacities for the global symmetries and we use the shorthand $\omega=\omega_1+\omega_2$.

\section{Pentagon identity for mirror symmetry}

For a special limit of flavor fugacities, the first duality  can be reduced to the mirror symmetry for $\mathcal N=2$ supersymmetric gauge theories. We will make the breaking of flavor symmetry on the level of partition functions. Here we follow the procedure described in \cite{Sarkissian:2022mlb}. First we take the limit $a_3 \to \infty$ and of course $b_3=\omega -a_1 - a_2 - a_3 - b_1 - b_2$ with the balancing condition $\sum_{i=1}^3 u_i + v_i = r$. If we collect terms coming from relevant Bernoulli polynomials which are not used in limit operations on the left-hand side, the following remains:
\begin{align} \nonumber
        \sum\limits_{y = 0}^{[r/2]} \epsilon(y)  \int_{-i\infty}^{i\infty} & e^{\frac{i\pi}{2} A } \prod _{i=1}^{2}\gamma^{(2)}(a_i -z +\omega_1(u_i - y); +\omega_1 r, \omega) \gamma^{(2)}(a_i - z) +\omega_2(r-(u_i - y)); \omega_2 r, \omega)\\ \nonumber
        &\times \gamma^{(2)}( b_i + z + \omega_1(v_i + y); \omega_1 r, \omega) \gamma^{(2)}(b_i + z + \omega_2(r-(v_i + y)); \omega_2 r, \omega) \frac{dz}{i r\sqrt{\omega_1\omega_2}}\\ \nonumber
        &= \prod\limits_{i,j=1}^{2} \gamma^{(2)}(a_i+ b_j +\omega_1(u_i + v_j); \omega_1 r, \omega)\gamma^{(2)}(a_i + b_j + \omega_2(r-(u_i + v_j)); \omega_2 r, \omega) \\ \nonumber
        & \qquad \quad \times \gamma^{(2)}(\omega - a_1 -b_1 -a_2-b_2 + \omega_1(r-u_1-v_1-u_2-v_2); \omega_1 r, \omega) \\
        & \qquad \quad \times \gamma^{(2)}(\omega - a_1 -b_1 -a_2-b_2 + \omega_2(u_1+v_1+u_2+v_2); \omega_2 r, \omega) ,
\end{align}
where the exponential factor is
\begin{equation*}
    \begin{split}
        A= \frac{1}{\omega_1 \omega_2 r} \big(2z(a_1+b_1+a_2+b_2) + (a_2+b_2)(-a_1+b_1)+(a_1+b_1)(-a_2+b_2)\big) \\
        + \frac{2y}{r}\big(u_1+v_1+u_2+v_2\big)- \frac{2}{\omega r}\big(\omega_1 - \omega_2 \big) \big(u_1 u_2 -v_1 v_2\big) .
    \end{split}
\end{equation*}
In the latter expression, we used the balancing condition to rewrite $u_3+v_3$ in terms of other flavor charges as $u_3+v_3=r-u_1-u_2-v_1-v_2$.

Now if we take $a_2 \to \infty$ limit, and define $a_1 + b_1 = g_1$, $a_2 +b_2=g_2$, $u_1 + v_1 = t_1$, $u_2 + v_2= t_2$ while sending\footnote{ Actually in the latter expression we only have the following combinations $u_1+v_1$ and $a_1+b_1$, therefore by sending we do not lose any information.} $b_1, v_1 \to 0$, we get:
\begin{align} \nonumber
        \sum\limits_{y = 0}^{[r/2]}  \int_{-i\infty}^{i\infty} &e^{\frac{2\pi i z}{\omega_1 \omega_2 r} (\frac{2g_1 + g_2 - \omega}{2}) + \frac{i \pi y}{r} (t_1  + 2t_2 - r)}\gamma^{(2)}(z +\omega_1 y; \omega_1 r, \omega) \gamma^{(2)}(z +\omega_2(r-y)); \omega_2 r, \omega)\\ \nonumber
        & \times \gamma^{(2)}(-z+g_1 +\omega_1(t_1 - y); \omega_1 r, \omega) \gamma^{(2)}(-z+g_1 +\omega_2(r-t_1+ y)); \omega_2 r, \omega)\frac{dz}{i r\sqrt{\omega_1\omega_2}} \\ \nonumber
        &=e^{\frac{2\pi i }{\omega_1 \omega_2 r} (\frac{g_1}{2}) (\frac{2g_2 + g_1 - \omega}{2}) + \frac{i \pi}{r} ( \frac{-t_1}{2})(t_1 + 2t_2 - r)} \gamma^{(2)}(\omega - g_1 - g_2 + \omega_1(r - t_1 -t_2); \omega_1 r, \omega)\\ \nonumber
        & \qquad \times \gamma^{(2)}(\omega - g_1 - g_2 + \omega_2(t_1 + t_2);\omega_2 r, \omega) \gamma^{(2)}(g_2 + \omega_1 t_2; \omega_1 r, \omega) \gamma^{(2)}( g_2 + \omega_2(r-t_2);\omega_2 r, \omega) \\
        & \qquad \times \gamma^{(2)}(g_1 + \omega_1 t_1; \omega_1 r, \omega)\gamma^{(2)}( g_1 + \omega_2(r- t_1);\omega_2 r, \omega)     \label{lastpentagon}
\end{align}
As one can see, the expression (\ref{lastpentagon}) has the form of the integral pentagon relation. The redefinition of the fugacities as $g_1/2 \to \alpha$, $z \to z + \alpha$, $g_2 \to \beta - \alpha $ and $t_1 \to 2n_{\alpha} $, $y \to y + n_{\alpha}$, $t_2 \to n_{\beta} - n_{\alpha}$ gives the integral identity (\ref{mirror}) for mirror symmetry \cite{Imamura:2012rq}. The case $r=1$ corresponds to the identity from \cite{Sarkissian:2022mlb,Faddeev:2000if}.

\section{Summary}

We report a solution to the pentagon equation obtained via reduction of equality for $S_b^3/\mathbb Z_r$ partition functions  of a certain $\mathcal N=2$ supersymmetric duality via the approach described in \cite{Sarkissian:2022mlb}.

There are many possible directions for future research.  The supersymmetric partition functions are related to topological invariants of triangulated three-manifolds via $3-3$ relation \cite{Dimofte:2011ju,Dimofte:2012pd}. The special  $r=1$ case of integral identities discussed here is used in \cite{Kashaev:2012cz} for the construction of the Turaev-Viro invariant of three-manifolds. It would be interesting to construct the invariant of the lens space \cite{suzuki2002turaev}. Much work remains to be done in this direction.

It would be interesting to construct a solution to the corresponding Yang-Baxter equation using the obtained solution to the pentagon equation in the context of gauge/YBE correspondence \cite{Gahramanov:2017ysd,Yamazaki:2018xbx}. 

Another possible future direction is obtaining the five-term identity \cite{Alexandrov:2015xir} implied by Kontsevich-Soibelman wall-crossing via the procedure described in \cite{Faddeev:2011tah}. 



\section*{Acknowledgements} 

It is pleasure to thank Mustafa Mullahasanoglu for the helpful discussions. The work of Ilmar Gahramanov is partially supported by TUBITAK grant 220N106, by the 1002-TUBITAK Quick Support Program under grant number 121F413, and by the Russian Science Foundation grant number 22-72-10122.

\bibliographystyle{utphys}
\bibliography{pentagon}

\end{document}